How would you solve the problem:
The height of a ball is measured by the function h(t)=-16t^2+112t-64, where h(t) is measured in feet and t is time in seconds.
(a) Evaluate h(2) and explain what dependent and independent variables represent.
(b) What is the height of the ball at 3 seconds?
(c) When is the ball at a height of 128 ft.?

Answer by Alan3354(69443)

How would you solve the problem:
The height of a ball is measured by the function h(t)=-16t^2+112t-64, where h(t) is measured in feet and t is time in seconds.
(a) Evaluate h(2) and explain what dependent and independent variables represent.
sub 2 for y
h(2) = -16*4 + 224 - 64 = 144 feet
h(2) is the height at 2 seconds, 2 is the time in seconds.
-----------------------
(b) What is the height of the ball at 3 seconds?
h(3) = -16*9 + 112*3 - 64 = 128 feet
---------------------
(c) When is the ball at a height of 128 ft.?
128 = -16t^2 + 112t - 64
t^2 - 7t - 12 = 0
t = 3 seconds (ascending)
t = 4 seconds (descending)